Let F n(x) =1/n *e^(-n^2*x^2) if x element of R ,n=1,2 .....

Prove that fn tend to 0 uniformly on R ,that derivative of fn tend to 0 point wise on R ,but that the convergence of {fn'} is not uniform on any interval containing the origin


1
Expert's answer
2021-03-10T10:07:14-0500

Here 

"\underset{{n\rightarrow \infty}}{lim }f_n(x)=0, \forall x"


And "f_n(x)=\dfrac{e^{-n^2x^2}}{n}"


"f'_n(x)=\dfrac{-2xn^2 e^{-n^2x^2}}{n}"


so "f'_n(x)=-2xne^{-n^2x^2}"



Now "-2xne^{-n^2x^2}" attains maximum value "\dfrac{2}{e} \text{ at } x=\dfrac{-1}{n}"


 tending to 0 as "n\rightarrow \infty."


Let us take the interval [a,b] containg 0.


Thus

  "M_n=\underset {x\in [a,b]}{Sup}|f_n(x)-f(x)|"


  "= \underset{x\in [a,b]}{Sup}|-2xne^{-n^2x^2}|"


     "=\dfrac{2}{e}" Which does not tends to zero as "n\rightarrow \infty."


 Hence the sequence "{f_n}" is not uniformly convergent in any interval [a,b] containing the origin.